OpenStudy (coming.home):

I'm having some more trouble with this algebraic expression. a=2, b=3, and c=-6 12 - 2(a-b)^2 So when you substitute the variables for numbers you get: 12 - 2 (2-3) ^2. I'm a little stuck after there. Can anyone help me?

OpenStudy (phi):

order of operations , do stuff in parens first what is 2-3 ?

OpenStudy (coming.home):

It would be -1.

OpenStudy (phi):

ok, now which operation do we do next ?

OpenStudy (coming.home):

Then you do exponents right? So the answer to that would be -1

OpenStudy (phi):

\[ 12 - 2 (2-3) ^2 \\ 12 - 2 (-1) ^2\] the exponent part is (-1)^2 which means -1*-1

OpenStudy (coming.home):

Ohhhh right. Okay so 1. The next step would be multiplication right? So that would be -2 * 1 which is... -2

OpenStudy (phi):

\[ 12 - 2 (2-3) ^2 \\ 12 - 2 (-1) ^2 \\ 12-2\cdot 1 \\ 12 -2 \]

OpenStudy (coming.home):

Awesome I got it right then! Thanks so much for your help,I understand now :) You were really helpful btw

OpenStudy (phi):

btw, we can treat 12 -2*1 as 12+ -2*1 which gives 12+ -2 = 10 or as 12 - 2*1 wich is 12 - 2 = 10 the first way we think of adding -2 then second way , we think of it as subtracting

OpenStudy (coming.home):

Oooooo thanks for the tip :)